When making a series of spectroscopic measurements where the concentration of a particular species is varied...

  1. When making a series of spectroscopic measurements where the concentration of a particular species is varied and measured using the same cell, why is it always good practice to start with the solution of lowest concentration and work your way up sequentially to the solution of highest concentration?
  2. With reference to the Stern-Volmer equation, state what the value of the vertical axis intercept is expected to be.
Answer #1
  • Generally we measure the concentration of a color solutions by spectroscopic measurement and we always try to measure lowest concentration to highest concentration. It is done to minimize the error to the concentration measurement. Here, we always use same cubic cell that contains the solution. At first, we measure the the lowest concentration and then we use the more concentrated solution because if a very small amount of previous solution is left in cell it does not effective change on the actual concentration to the next solution. But if we first measure the highest concentration then we measure the lower concentration a very little amount of previous solution can happen a effective change to the concentration of the next solution. So, we always try to measure lowest concentration to highest concentration.
  • The Stern-Volmer equation is

I0/I = 1 + Ksv[Q]

I0= absorbance of the solution when quencher concentration is zero.

I= absorbance of the solution when quencher concentration

is [Q].

Ksv= Stern-Volmer constant.

This Stern-Volmer equation is like a y=mx+c type equation where c is a intercept along y axis or vertical axis.

So, if we plot I0/I vs [Q] we will get a intercept along y axis or vertical axis and according to Stern-Volmer equation the value of intercept is 1.